OpenStudy (anonymous):

Calculus: Find the derivative. Given the equation is s(t)=t^4-16t^2+5 For what values of t is v(t)=0

OpenStudy (anonymous):

\[t ^{4}-16t ^{2}+5\]

OpenStudy (anonymous):

\[4t ^{3}-32t\]

OpenStudy (anonymous):

Diff of s(t) is v(t)

OpenStudy (anonymous):

v(t)=ds/dt=4t^3-32t=0 thus, \[4t ^{3}-32t=0\] \[4t(t ^{2}-32)=0\]

OpenStudy (anonymous):

sorry last one is \[4t(t ^{2}-8)=0\]

OpenStudy (anonymous):

either t=0 or t^2-8=0 either t=0 or t^2=8 either t=0 or t=+sqrt(2) or t=-sqrt(2)

OpenStudy (anonymous):

wait i got thx

OpenStudy (anonymous):

welcome.......